OpenStudy (anonymous):

Solve 5x2 = –30x – 65

OpenStudy (anonymous):

x=-6+4i/2,-6-4i/2

OpenStudy (anonymous):

how did you come up with that answer

OpenStudy (anonymous):

i use an online calculator

OpenStudy (anonymous):

@Insertcoolmetephore may i get the website to this calculator

OpenStudy (anonymous):

i only use this site when i forget how to do a problem. Mathaway

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

\(5x^2=–30x–65\\5x^2+30x+65=0\\x^2+6x+13=0\\x=\frac{-6\pm\sqrt{6^2-4(13)}}{2}=-3\pm\sqrt{-4}=-3\pm2i\)

OpenStudy (anonymous):

Don't use a website to do the problem, use it to check your work... all that's needed to do this problem is the quadratic formula.

OpenStudy (anonymous):

that too